test ∑n=1 ∞​ [✓n^4 +9 - ✓n^4 -9] is convergent or not


1
Expert's answer
2021-06-29T09:29:18-0400

"\Sigma_{n=1}^{\infty}(\sqrt{n^4+9}-\sqrt{n^4-9})=\\
=\Sigma_{n=1}^{\infty}\frac{(\sqrt{n^4+9}-\sqrt{n^4-9})(\sqrt{n^4+9}+\sqrt{n^4-9})}{(\sqrt{n^4+9}+\sqrt{n^4-9})}=\\
=\Sigma_{n=1}^{\infty}\frac{18}{\sqrt{n^4+9}+\sqrt{n^4-9}}\\
\frac{18}{\sqrt{n^4+9}+\sqrt{n^4-9}}\leftrightarrow\frac{18}{2n^2}=\frac{9}{n^2}"

"\Sigma_{n=1}^{\infty}\frac{9}{n^2}" is convergent


"\Sigma_{n=1}^{\infty}(\sqrt{n^4+9}-\sqrt{n^4-9})" is convergent